# Orientation reconstruction of transformation titanium alloys via polarized light microscopy: methodology and assessment

### v1

@@ 4 Polycrystal Demonstration
To demonstrate, we will utilize a virtual polycrystalline sample in which we generate a to transformation microstructure. This enables us to validate our results against the known (true) deterministic orientations, in which the field of both the room temperature colonies and the parent grains are both fully and unambiguously known.
@@
Next, we employ the method described in Section 3.2 to calculate the two possible parent grain orientations for each of the sets of clustered colonies. We note that of the two sets back-calculated orientations, one of the sets is the true parent grain orientations (presented previously in Figure 3). We further present the second set of reflected parent grain orientations that we would otherwise be unable to distinguish between in Figure 5. From these two parent grain orientations, we finally calculate the 24 (total) variants (Section 3.3), and compare their axes against each colony’s measured axis and its reflection. In total, four variants are possible for each measured fiber, and are presented in Figure 3 (i.e., the true orientation set), as well as the three other possible orientation sets in Figure 6. In total, we demonstrate that the fibers measured via PLM allow for the determination of the colony orientations to any of the possibilities presented in Figures 3 and 6.
@@ 5 Crystal Plasticity Simulations
In a previous study (van Wees et al., 2023), we demonstrated that the random choice of the third Euler-Bunge angle (i.e., based on what PLM was able to measure) had a potentially large effect on the predicted deformation response of crystal plasticity finite element (CPFEM) simulations. Specifically, we found that the elastic responses between samples in which the third Euler-Bunge angle was randomly chosen did not differ appreciably, as expected owing to the transversely isotropic nature of HCP crystals (Bower, 2009; Nye, 1957) and thus the negligible impact of rotation about the axis on purely elastic response. However, as plasticity developed, we observed that the deviations in stress predictions between samples grew considerably to average errors of approximately 20%, to the point where PLM could not be effectively used as an instantiation method for CPFEM simulations in which the prediction past the elastic regime was of interest.
@@ 5.1.1 Microstructure Morphology
To generate idealized microstructures for use in simulations, we utilize multilevel tessellations (first described in (Kasemer et al., 2017a) for similar use) as realized by the software package Neper (Quey et al., 2011). In multilevel tessellations, we may perform successive levels of discretization by considering the sub-volumes from the previous scale of tessellation as domains available for further discretization, overall in an effort to reach a desired morphology. Here, we aim to maintain the parent grain morphology with sub-grain discretization to create volumes representing colonies which collectively maintain their parent grain morphology.
@@ 5.1.2 Orientation Generation
Regarding the selection of the parent grain orientations, specifically, we assume two primary paradigms. In the first, we generate a random set of parent grain orientations. In the second, we generate a set of parent grain orientations assuming a cube texture (Raabe et al., 2005), as Ti64 is often processed (rolled) above the transus temperature when it is entirely phase (BCC), and cubic materials tend toward a cube texture during rolling. We plot the parent grains with three different texture assumptions (random, and two different strengths of cube texture) in Figure 8. We then plot the ultimate phase microstructures that arise from these parent grains in Figure 9.
@@ 5.1.3 Orientation Variation
As described in Section 5.1.2, we ultimately have three base samples (as depicted in Figure 9). To understand the effect that the quasi-determinsitic approach to identifying the colony orientations (described in Section 3) has on the predicted deformation response, we generate a set of permuted samples in which the colony orientations are randomly chosen among the four possibilities that each colony may be based on the quasi-deterministic reconstruction method described in this study. This mimics the degree to which we can determine the colony orientations via our quasi-deterministic reconstruction method. We generate 10 different samples for each of the three base microstructures, where each of the colony orientations are randomly chosen among their four possibilities (i.e., 30 permuted samples in total). The predicted deformation response of these permuted samples are compared against the predicted deformation response of the base samples (i.e., the true samples). Further, we will compare the results of the simulations described here to those from a previous study (van Wees et al., 2023) in which the third Euler-Bunge angle was chosen entirely at random to demonstrate the superior predictive capability when utilizing orientations from the reconstruction method.
@@ 5.2 Material and Modeling Parameters
We utilize the elastic and plastic modeling parameters for the phase of Ti-6Al-4V (Ti64) in this study. We choose Ti64, as it is exhibits a transformation microstructure that adheres to the Burgers orientation relationship (Lütjering and Williams, 2003). Further, Ti64 has enjoyed a significant degree of study (Kasemer et al., 2017a, b; Wielewski et al., 2017; Dawson et al., 2018), and consequently we have a high degree of confidence in the elastic and plastic modeling parameters. We summarize the crystal plasticity modeling parameters in Tables 1 and 2. We do not consider the retained phase morphology in this study, as it represents a small volume fraction of the material at room temperature (less than 10% total volume), is often segregated to grain boundaries (as in the case of the microstructure realized by the mill annealed processing route (Lütjering and Williams, 2003; Kasemer et al., 2017a)), and due to the relatively high degree of uncertainty in its modeling parameters. Consequently, we construct our virtual samples considering only the phase (i.e., single phase simulations).
| 169.7 | 88.7 | 61.7 | 42.5 |
| 390 | 468 | 663 | 0.01 | 1 | 190 | 530 |
@@ 5.3 Boundary Conditions and Load History
We deform the specimens in uniaxial tension at a rate of (i.e., quasi-static) in the sample direction using minimal boundary conditions (Kasemer, 2024) that minimize boundary conditions effects not representative of uniaxial tension. We apply deformation through 50 equally spaced time steps to a strain of (i.e., 50 time steps of ), which allows for the close inspection of various points along the loading history.
@@ 5.4 Simulation Results
Here, we present the results of the simulations. To gain a broad understanding of the effects of the choice of orientations on the predictions of the deformation field, we compare the equivalent stress fields between the three base sample and their permutations with randomly varied orientations. We borrow the metric utilized in (van Wees et al., 2023) to quantify the error in stress predictions element-to-element between a single permuted sample and the base sample:
@@
We observe that, generally, the results of the permuted samples compare more favorably to the base samples in the elastic regime compared to in the plastic regime. At macroscopic strain states of and (the latter shortly before macroscopic yield), the average errors across all three samples are 2.04% and 2.53%, respectively, while at macroscopic strain states of and (i.e., after macroscopic yield), the average errors across are three samples are 5.26% and 5.47%, respectively. The lower errors in the nominally elastic regime are consistent with the elastic transverse isotropic nature of HCP crystals (Bower, 2009), where we would expect rotations about the axes of crystals to have limited effect on the predicted stress field (rotations of the axes due to reflections, however, could have stronger effects). As plasticity develops, the differences in slip system activity will lead to more pronounced differences in the predicted deformation fields.
We next compare to our previous study (van Wees et al., 2023) in which the third Euler-Bunge angle was chosen entirely at random without employing the quasi-deterministic reconstruction method presented here. While results in the elastic regime between the two studies are broadly comparable (again, a consequence of the elastic transverse isotropic nature of HCP crystals), the results in the plastic regime differ greatly. Specifically, when utilizing a random choice of the four possible orientations as we have described in this study, the permuted simulations exhibit a significant reduction in error compared to when the third Euler-Bunge angle is chosen entirely at random. At , for example, the average error in stress fields compared to the base samples across all three microstructures when using our quasi-deterministic reconstruction method is 5.47%, compared to average deviations of approximately 23% in our previous study when using entirely randomly chosen values for the third Euler-Bunge angle. In other words, randomly choosing one of the four quasi-deterministic orientations results in a more than four-fold reduction in deviation of stress field predictions from truth compared to an entirely random choice of the third Euler-Bunge angle. This large reduction in uncertainty lends significant credibility to PLM as a characterization method for informing CPFEM simulations of titanium alloys.
@@ 6 Conclusion
In this study, we have presented a method to reconstruct quasi-deterministic orientations for titanium measured via polarize light microscopy. Our method exploits the Burgers orientation relationship (BOR) to back-calculate the likely orientations of the high temperature parent grains, from which the BOR can be used to forward-calculate the deterministic room-temperature colony orientations. We can match these ideal variants against the PLM-measured fibers, and show that based on the ambiguities present in PLM, we can narrow down the orientations to four possibilities. We demonstrated our reconstruction method on a synthetic sample containing fully-known orientations for both the room temperature colony field as well as the parent grain field, which provides validation for our method’s ability to calculate the quasi-deterministic orientations.
@@ Appendix A Demonstration of Reconstruction Method in Rodrigues Space
Here, we provide a demonstration of our reconstruction method in Rodrigues orientation space. The Rodrigues orientation parameterization (Frank, 1988) is advantageous as it allows for the representation of all unique orientations via a relatively simple closed fundamental region, and due to the fact that crystallographic fibers present as straight lines in Rodrigues space (Kumar and Dawson, 1998; Quey et al., 2015). To demonstrate this property, we present a visualization of three random orientations plotted in the hexagonal symmetry fundamental region of Rodrigues space in Figure A.1(a), along with their respective fibers of orientations who share the same axis orientations plotted as lines. PLM alone is unable to naturally distinguish between any of the orientations that lie on each line, since all orientations on each line share the same axis orientation (i.e., they are axis similar). Thus, this figure demonstrates the consequences of PLM’s first limitation. We note that due to crystal symmetry, a fiber may exit one side of the fundamental region and appear on the opposite side, and thus all orientations which share the same axis may not be on a single line (though are collectively of the same fiber) (Kumar and Dawson, 1998). Further, we plot the reflected orientations and fibers in Figure A.1(b).
@@ Appendix B Crystal Plasticity Finite Element Method
In this study, we utilize the CPFEM software package FEPX (Kasemer, 2024; Quey and Kasemer, 2022). FEPX is a free and open-source CPFEM solver which is MPI parallelized, allowing for its deployment on large-scale computational clusters. FEPX is thus capable of high-throughput simulations of the elastic-plastic deformation response of high-fidelity representations of polycrystalline microstructures. FEPX assumes quasi-static, ductile, isothermal deformation, and consequently the models it employs reflects these assumptions. As FEPX uses widely-accepted models employed in a well-established solution approximation scheme (i.e., a non-linear finite element solver), we present here a truncated description of the most important portions of the model for brevity. We refer the reader to (Dawson and Boyce, 2015) for a full description of the kinematics, models, and finite element implementation. The models below are considered locally at points within each finite element in a polycrystalline mesh.
@@
where C is the fourth-order anisotropic stiffness tensor. The stiffness tensor reflects the symmetry of the crystal (Bower, 2009; Nye, 1957), and is orientation dependent, written above as a function of the orientation parameterized as a Rodrigues vector, .
For modeling plastic deformation, we consider slip kinetics via a rate-dependent restricted-slip power law (Nemat-Nasser and Obata, 1986) model:
@@
Here, is the lattice spin vector, described in full in (Dawson and Boyce, 2015), itself a function of the amount of local plastic plastic shearing.

### v2

@@ 4 Polycrystal Demonstration
To demonstrate, we will utilize a virtual polycrystalline sample in which we generate a to transformation microstructure. While we envision that this method can and will be applied to experimental data, we opt to focus this study on the establishment of the method via the inspection of the idealized problem—i.e., one in which the sample contains colonies of only a single orientation, where the transformation is perfect, and there exists multiple colony variants within each parent grain. We note that while similar parent grain orientation fields may be reconstructed from EBSD data using commercial EBSD software, these fields are inferred rather than directly experimentally observed, and thus comparison against these results—which are like any other experimental data prone to uncertainty and sensitivity to choice of reconstruction parameters—is subpar when there exists an alternative that is perfect and unambiguous. The construction of virtual samples enables us to deterministically validate our method and software against known (true) deterministic orientations in which the field of both the room temperature colonies and the parent grains are both fully and unambiguously known.
@@
Next, we employ the method described in Section 3.2 to calculate the two possible parent grain orientations for each of the sets of clustered colonies. We note that of the two sets back-calculated orientations, one of the sets is the true parent grain orientations (presented previously in Figure 3). We further present the second set of reflected parent grain orientations that we would otherwise be unable to distinguish between in Figure 5. We take care here to note that our method is unable to distinguish between the orientations presented in Figure 3 and Figure 5 independently per grain—i.e., the reader should not misconstrue these figures as the only two possible fields of orientations, and indeed each grain each can take one of two possible orientations. From these two parent grain orientations, we finally calculate the 24 (total) variants (Section 3.3), and compare their axes against each colony’s measured axis and its reflection. In total, four variants are possible for each measured fiber, and are presented in Figure 3 (i.e., the true orientation set), as well as the three other possible orientation sets in Figure 6. We again take care here to note that our method is unable to distinguish between the orientations presented in Figure 3 and Figure 6 independently per colony—i.e., the reader should not misconstrue these figures as the only four possible fields of orientations, and indeed each colony each can take one of four possible orientations. In total, we demonstrate that the fibers measured via PLM allow for the determination of the colony orientations to any of the possibilities presented in Figures 3 and 6.
@@ 5 Crystal Plasticity Simulations
In a previous study [28], we demonstrated that the random choice of the third Euler-Bunge angle (i.e., based on what PLM was able to measure) had a potentially large effect on the predicted deformation response of crystal plasticity finite element (CPFEM) simulations. Specifically, we found that the elastic responses between samples in which the third Euler-Bunge angle was randomly chosen did not differ appreciably, as expected owing to the transversely isotropic nature of HCP crystals [36, 37] and thus the negligible impact of rotation about the axis on purely elastic response. However, as plasticity developed, we observed that the deviations in stress predictions between samples grew considerably to average errors of approximately 20%, to the point where PLM could not be effectively used as an instantiation method for CPFEM simulations in which the prediction past the elastic regime was of interest.
@@ 5.1.1 Microstructure Morphology
To generate idealized microstructures for use in simulations, we utilize multilevel tessellations (first described in [32] for similar use) as realized by the software package Neper [38]. In multilevel tessellations, we may perform successive levels of discretization by considering the sub-volumes from the previous scale of tessellation as domains available for further discretization, overall in an effort to reach a desired morphology. Here, we aim to maintain the parent grain morphology with sub-grain discretization to create volumes representing colonies which collectively maintain their parent grain morphology.
@@ 5.1.2 Orientation Generation
Regarding the selection of the parent grain orientations, specifically, we assume two primary paradigms. In the first, we generate a random set of parent grain orientations. In the second, we generate a set of parent grain orientations assuming a cube texture [41], as Ti64 is often processed (rolled) above the transus temperature when it is entirely phase (BCC), and cubic materials tend toward a cube texture during rolling. We plot the parent grains with three different texture assumptions (random, and two different strengths of cube texture) in Figure 8. We then plot the ultimate phase microstructures that arise from these parent grains in Figure 9.
@@ 5.1.3 Orientation Variation
As described in Section 5.1.2, we ultimately have three base samples (as depicted in Figure 9). To understand the effect that the quasi-determinsitic approach to identifying the colony orientations (described in Section 3) has on the predicted deformation response, we generate a set of permuted samples in which the colony orientations are randomly chosen among the four possibilities that each colony may be based on the quasi-deterministic reconstruction method described in this study. This mimics the degree to which we can determine the colony orientations via our quasi-deterministic reconstruction method. We generate 10 different samples for each of the three base microstructures, where each of the colony orientations are randomly chosen among their four possibilities (i.e., 30 permuted samples in total). The predicted deformation response of these permuted samples are compared against the predicted deformation response of the base samples (i.e., the true samples). Further, we will compare the results of the simulations described here to those from a previous study [28] in which the third Euler-Bunge angle was chosen entirely at random to demonstrate the superior predictive capability when utilizing orientations from the reconstruction method.
@@ 5.2 Material and Modeling Parameters
We utilize the elastic and plastic modeling parameters for the phase of Ti-6Al-4V (Ti64) in this study. We choose Ti64, as it is exhibits a transformation microstructure that adheres to the Burgers orientation relationship [30]. Further, Ti64 has enjoyed a significant degree of study [32, 42, 43, 44], and consequently we have a high degree of confidence in the elastic and plastic modeling parameters. We summarize the crystal plasticity modeling parameters in Tables 1 and 2. We do not consider the retained phase morphology in this study, as it represents a small volume fraction of the material at room temperature (less than 10% total volume), is often segregated to grain boundaries (as in the case of the microstructure realized by the mill annealed processing route [30, 32]), and due to the relatively high degree of uncertainty in its modeling parameters. Consequently, we construct our virtual samples considering only the phase (i.e., single phase simulations).
169.7 | 88.7 | 61.7 | 42.5 |
390 | 468 | 663 | 0.01 | 1 | 190 | 530 |
@@ 5.3 Boundary Conditions and Load History
We deform the specimens in uniaxial tension at a rate of (i.e., quasi-static) in the sample direction using minimal boundary conditions [45] that minimize boundary conditions effects not representative of uniaxial tension. We apply deformation through 50 equally spaced time steps to a strain of (i.e., 50 time steps of ), which allows for the close inspection of various points along the loading history.
@@ 5.4 Simulation Results
Here, we present the results of the simulations. To gain a broad understanding of the effects of the choice of orientations on the predictions of the deformation field, we compare the equivalent stress fields between the three base sample and their permutations with randomly varied orientations. We borrow the metric utilized in [28] to quantify the error in stress predictions element-to-element between a single permuted sample and the base sample:
@@
We observe that, generally, the results of the permuted samples compare more favorably to the base samples in the elastic regime compared to in the plastic regime. At macroscopic strain states of and (the latter shortly before macroscopic yield), the average errors across all three samples are 2.04% and 2.53%, respectively, while at macroscopic strain states of and (i.e., after macroscopic yield), the average errors across are three samples are 5.26% and 5.47%, respectively. The lower errors in the nominally elastic regime are consistent with the elastic transverse isotropic nature of HCP crystals [36], where we would expect rotations about the axes of crystals to have limited effect on the predicted stress field (rotations of the axes due to reflections, however, could have stronger effects). As plasticity develops, the differences in slip system activity will lead to more pronounced differences in the predicted deformation fields.
We next compare to our previous study [28] in which the third Euler-Bunge angle was chosen entirely at random without employing the quasi-deterministic reconstruction method presented here. While results in the elastic regime between the two studies are broadly comparable (again, a consequence of the elastic transverse isotropic nature of HCP crystals), the results in the plastic regime differ greatly. Specifically, when utilizing a random choice of the four possible orientations as we have described in this study, the permuted simulations exhibit a significant reduction in error compared to when the third Euler-Bunge angle is chosen entirely at random. At , for example, the average error in stress fields compared to the base samples across all three microstructures when using our quasi-deterministic reconstruction method is 5.47%, compared to average deviations of approximately 23% in our previous study when using entirely randomly chosen values for the third Euler-Bunge angle. In other words, randomly choosing one of the four quasi-deterministic orientations results in a more than four-fold reduction in deviation of stress field predictions from truth compared to an entirely random choice of the third Euler-Bunge angle. This large reduction in uncertainty lends significant credibility to PLM as a characterization method for informing CPFEM simulations of titanium alloys.
@@ 6 Conclusion
In this study, we have presented a method to reconstruct quasi-deterministic orientations for titanium measured via polarize light microscopy. Our method exploits the Burgers orientation relationship to back-calculate the likely orientations of the high temperature parent grains, from which the Burgers orientation relationship can be used to forward-calculate the deterministic room-temperature colony orientations. We can match these ideal variants against the PLM-measured fibers, and show that based on the ambiguities present in PLM, we can narrow down the orientations to four possibilities. We demonstrated our reconstruction method on a synthetic sample containing fully-known orientations for both the room temperature colony field as well as the parent grain field, which provides validation for our method’s ability to calculate the quasi-deterministic orientations.
@@ Appendix A Demonstration of Reconstruction Method in Rodrigues Space
Here, we provide a demonstration of our reconstruction method in Rodrigues orientation space. The Rodrigues orientation parameterization [46] is advantageous as it allows for the representation of all unique orientations via a relatively simple closed fundamental region, and due to the fact that crystallographic fibers present as straight lines in Rodrigues space [35, 47]. To demonstrate this property, we present a visualization of three random orientations plotted in the hexagonal symmetry fundamental region of Rodrigues space in Figure A.1(a), along with their respective fibers of orientations who share the same axis orientations plotted as lines. PLM alone is unable to naturally distinguish between any of the orientations that lie on each line, since all orientations on each line share the same axis orientation (i.e., they are axis similar). Thus, this figure demonstrates the consequences of PLM’s first limitation. We note that due to crystal symmetry, a fiber may exit one side of the fundamental region and appear on the opposite side, and thus all orientations which share the same axis may not be on a single line (though are collectively of the same fiber) [35]. Further, we plot the reflected orientations and fibers in Figure A.1(b).
@@ Appendix B Crystal Plasticity Finite Element Method
In this study, we utilize the CPFEM software package FEPX [45, 40]. FEPX is a free and open-source CPFEM solver which is MPI parallelized, allowing for its deployment on large-scale computational clusters. FEPX is thus capable of high-throughput simulations of the elastic-plastic deformation response of high-fidelity representations of polycrystalline microstructures. FEPX assumes quasi-static, ductile, isothermal deformation, and consequently the models it employs reflects these assumptions. As FEPX uses widely-accepted models employed in a well-established solution approximation scheme (i.e., a non-linear finite element solver), we present here a truncated description of the most important portions of the model for brevity. We refer the reader to [48] for a full description of the kinematics, models, and finite element implementation. The models below are considered locally at points within each finite element in a polycrystalline mesh.
@@
where C is the fourth-order anisotropic stiffness tensor. The stiffness tensor reflects the symmetry of the crystal [36, 37], and is orientation dependent, written above as a function of the orientation parameterized as a Rodrigues vector, .
For modeling plastic deformation, we consider slip kinetics via a rate-dependent restricted-slip power law [5] model:
@@
Here, is the lattice spin vector, described in full in [48], itself a function of the amount of local plastic plastic shearing.